OpenStudy (anonymous):

write expression as sine cosine or tangent of single angle sin 48° cos 15° -cos48° sin15° open study

zepdrix (zepdrix):

Hey Hannah, apply your Angle Difference Formula for Sine:\[\large\rm \sin A \cos B-\cos A \sin B\quad=\quad \sin\left(A-B\right)\]